# FastTrackTr:Towards Real-time Multi-Object Tracking with Transformers

### v3

@@ Abstract
Transformer-based multi-object tracking (MOT) methods have captured the attention of many researchers in recent years. However, these models often suffer from slow inference speeds due to their structure or other issues. To address this problem, we revisited the Joint Detection and Tracking (JDT) method by looking back at past approaches. By integrating the original JDT approach with some advanced theories, this paper employs an efficient method of information transfer between frames on the DETR, constructing a fast and novel JDT-type MOT framework: FastTrackTr. Thanks to the superiority of this information transfer method, our approach not only reduces the number of queries required during tracking but also avoids the excessive introduction of network structures, ensuring model simplicity. Experimental results indicate that our method has the potential to achieve real-time tracking and exhibits competitive tracking accuracy across multiple datasets.
Figure 1: A comparison of the FPS, HOTA and MOTA of various versions of FastTrackTr and the most advanced transformer-based models on a single NVIDIA 4090. For the left image, the horizontal axis is FPS, the vertical axis is HOTA, and the radius of the shape is MOTA.
@@ 1 Introduction
Multi-object tracking (MOT) has long been a computer vision task of great interest to researchers and developers, with extensive application prospects across various industries. Currently, the mainstream tracking models are primarily online models, and the paradigm that has garnered widespread attention in recent years within online models remains tracking-by-detection (TBD). This paradigm, originating from SORT [2], has stood the test of time and demonstrates remarkable vitality. Its improved derivative methods [21, 8, 43, 1] continue to be considered state-of-the-art (SOTA).
In recent years, some tracking methods based on transformer [36] have also attracted considerable attention. However, these methods [33, 22, 45, 41, 11, 44, 12] face a significant challenge: suboptimal inference speed. There are various reasons for this, such as the slow performance of the MOTR series, which relies on query-based tracking. The main reason is not the network structure, but issues related to hardware acceleration caused by the variable number of queries (detailed justification can be found in the appendix). Both our experiments and other works [24] have also demonstrated that the inference speed after lightening the network is still not very ideal. Regarding another recent method, MOTIP [12], while it addresses many of MOTR’s issues, its main problem lies in its complex network structure. Even after replacing MOTIP’s Deformable DETR [53] with RT-DETR [50], its inference speed remains unsatisfactory (results shown in Figure 1). And, during its implementation, some of the data is dynamic, which makes it relatively difficult to accelerate using TensorRT.
Methods following the tracking-by-detection paradigm often claim to be extremely fast, but in reality, aside from those purely based on kinematics [4], the speed of most MOT systems is not ideal once the time required for the detection model is considered. Purely kinematic models typically perform worse than when they are combined with reid networks. So, is there a way to achieve both fast and accurate tracking? To address this question, we must look back at the JDT (Joint Detection and Tracking) paradigm, which was initially renowned for its speed. Although it has fallen out of favor in recent years due to various reasons, it likely represents a solution to the aforementioned issues. Previously, there was a transformer-based JDT model, known as the TransTrack [33], which shares the same pain points as MOTIP, including having two decoders. Other methods [22, 40, 3], similar to MOTR, track based on queries. CNN-based JDT models [47, 37], on the other hand, typically produce detection results and their corresponding id embeddings for association matching, with their main issue being that the models are outdated, hence they do not keep pace with current standards.
@@
Based on the above discussion, we present FastTrackTr, a Fast multi-object Tracking method grounded in the Transformer architecture and aligned with the JDT paradigm. This tracking method achieves impressive tracking progress through a unique decoder that integrates past information. Additionally, due to the ease of generating masks for this method, the batch size can be set relatively large, making our model considerably more training-friendly compared to some of the latest transformer-based MOT methods. Additionally, due to the ease of generating masks for this method, the batch size can be set relatively large, making our model considerably more training-friendly compared to some of the latest transformer-based MOT methods.
Furthermore, the deterministic data architecture inherent in FastTrackTr enables seamless TensorRT acceleration, achieving genuine real-time performance. This computational efficiency remains largely unattainable for contemporary transformer-based MOT methods due to their inherent dynamic computation patterns.
@@
    Extensive experiments on multiple benchmark datasets, including DanceTrack [34], SportsMOT [7] and MOT17 [23], demonstrate that our model not only achieves competitive accuracy SOTA methods, but also significantly improves inference speed.
@@ 2.1 Transformer-based MOT Methods
The earliest Transformer-based MOT methods were TransTrack [33] and TrackFormer [22]. The former is a JDT model, while the latter introduced the concept of track queries, utilizing these special queries to achieve tracking.
Subsequently, MOTR [45] proposed tracklet-aware label assignment (TALA), which optimized label assignment during the training phase to achieve end-to-end tracking. Later improvements to MOTR, such as MOTRv2 [49], MeMOTR [11], MOTRv3 [44], and CO-MOT [41], demonstrated the powerful tracking performance of the MOTR paradigm. Some efforts have been made to address MOTR’s slow inference speed through model lightweighting, but the results have been less than ideal. Recently, alternative Transformer architectures have been proposed. MOTIP’s [12] novel approach of directly predicting tracked object ID numbers is refreshing, while PuTR’s [16] use of Transformers to directly construct a tracking-by-detection paradigm tracker has also been inspirational.
@@ 2.2 Tracking-by-Detection
The TBD paradigm can be considered one of the simplest and most practical tracking methods. Since it essentially associates detected targets, this class of methods often yields satisfactory tracking accuracy as long as the detection results are not poor. The initial SORT algorithm simply utilized the motion state of detected objects for association. Subsequent models like DeepSORT [38] and ByteTrack [48] improved tracking accuracy by enhancing detection precision and employing embedding-based Re-identification (Re-id) networks. Some approaches, like OC-SORT [4] and HybridSORT [43], improved upon SORT using purely kinematics-based methods. Later models such as MotionTrack [26], DeepOC-SORT [21], and SparseTrack [18] demonstrated that simultaneous improvements in both Re-id network utilization and kinematic models yield the most significant performance enhancements. Currently, there are also algorithms that directly use networks for association. Besides the aforementioned PuTR [16], SUSHI [6] is a tracking method based on graph neural networks.
@@ 2.3 Joint Detection and Tracking
Joint detection and tracking aims to achieve detection and tracking simultaneously in a single stage. This approach was common a few years ago but has been seen less frequently in recent years. Here, we briefly review its development history. D&T [9] first proposed a multi-task architecture based on frame-based object detection and cross-frame trajectory regression. Subsequently, Integrated-Detection [54] improved detection performance by combining detection bounding boxes in the current frame with trajectories from the previous frame. Tracktor [32] directly used tracking boxes from the previous frame as candidate regions, then applied bounding box regression to provide tracking boxes for the current step, thus eliminating the box association process. JDE [37] and FairMOT [47] learned object detection and appearance embedding tasks from a shared backbone network. CenterTrack [51] localized targets through tracking-conditioned detection and predicted their offsets from the previous frame. TransTrack used two decoders to handle tracking and detection separately, then matched and associated the results from both decoders. Our FastTrackTr, by introducing a novel decoder, returns to the essence of the JDT method, simplifies the network structure, and achieves fast tracking.
@@ 3.1 FastTrackTr Architecture
Figure 2: The complete structure of FastTrackTr. For the initial frame, the model functions as a detection system, assigning unique identifiers to each detected object. Subsequently, from the second frame onwards, the historical queries from previous frames and their corresponding masks undergo processing through a single-layer encoder. The output and image feature of this frame are then fed into a historical-cross decoder. Based on the head’s output, detection results and ID embeddings are generated. These are then matched and associated with previous results to produce the final tracking outcome. This architecture achieves the efficient tracking, which can be run in real time on some devices.
The overall structure of FastTrackTr is illustrated in Figure 2. In frame $T_{1}$, a standard decoder is used to initialize the initial historical information. Subsequently, in the following frames, a Historical Encoder is employed to initialize this historical-representative information. In the next frame, a Historical Decoder, which includes a historical cross-attention layer compared to normal decoder layers, processes this historical information. The tracking principle is very similar to that of JDT and FairMOT, utilizing an additional ID Embedding head to obtain the object’s appearance embedding, which is then used along with detection results for association and matching. It is worth noting that the backbone network is ResNet [14].
@@ 3.2 Learning Detection and ID Embedding
To enable our FastTrackTr to simultaneously output both the target’s position and appearance embedding in a single forward pass, we formulate the problem as follows: Suppose we have a training dataset ${\mathbf{I},\mathbf{B},\mathbf{y}}_{i=1}^{N}$. Here, $\mathbf{I}\in\mathbb{R}^{c\times h\times w}$ represents an image frame, and $\mathbf{B}\in\mathbb{R}^{k\times 4}$ denotes the bounding box annotations for the $k$ targets in this frame. $\mathbf{y}\in\mathbb{Z}^{k}$ represents the partially annotated identity labels, where -1 indicates targets without an identity label. We hope model can output predicted bounding boxes $\hat{\mathbf{B}}\in\mathbb{R}^{\hat{k}\times 4}$ and appearance embeddings $\hat{\mathbf{F}}\in\mathbb{R}^{\hat{k}\times D}$, where $D$ is the dimension of the embedding.
For the output box, this requirement is easily met, as DETR[5] itself is a detection model. To achieve rapid performance, this paper employs RT-DETR [50] as the base model, which inherently possesses state-of-the-art detection capabilities. We have not made significant modifications to its detection component. The loss function for this part remains consistent with RT-DETR, as shown below:
[…]
where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}=$$\{\hat{c},\hat{b}\}$ and $y=\{c,b\},c$ and $b$ represent categories and bounding boxes, respectively. Similar to RT-DETR, we incorporate the IoU score into the objective function of the classification branch, following the approach from VFL [46], to enforce consistency between the classification and localization of positive samples. $\mathcal{L}_{box}$ includes the L1 loss and the generalized IoU loss [27], which are not described here. And, the auxiliary loss from DAB-DETR [17] is also employed in our approach.
The second objective addresses a metric learning problem, specifically the learning of an embedding space where instances of the same identity are close to each other while those of different identities are far apart. Previous JDT methods (e.g., JDE [37], FairMOT [47]) rely on one-hot labels to construct ID embeddings, necessitating predefined ID quantities to configure embedding dimensions. This design introduces two critical limitations: (1) explosive dimensionality growth in dense scenes and (2) severely restricted generalization due to dataset-specific ID dependencies. Subsequent Transformer-based trackers inherit these constraints through their query mechanisms. While recent approaches like MOTIP [12] (using ID dictionaries) and PuTR (leveraging embedding similarity) improve generalizability, their trajectory decoders overemphasize motion patterns, conflicting with FastTrackTr’s queries that encode richer spatial features and positional priors. This mismatch hinders effective convergence. To resolve these issues, we reformulate ID embedding learning as a re-identification (ReID) problem, decoupling embedding dimensions from ID quantities and prioritizing feature distance/similarity optimization. This paradigm shift enhances cross-scene adaptability and better aligns with image-driven tracking frameworks.
@@
For the above reasons and for better performance, in this paper, we adopt the more advanced circle loss [35] in this work. Specifically, by utilizing the dot product similarity, the id embeddings of the same object across multiple frames are regarded as positive pairs indicated as $s_{p}$, while those between distinct objects are treated as negative samples represented by $s_{n}$. - $\alpha_{p}$ and $\alpha_{n}$ are the adaptive weighting factors for positive and negative samples, determined by the following equations:
$\alpha_{p}=\left[O_{p}-s_{p}\right]^{+},\quad\alpha_{n}=\left[s_{n}-O_{n}\right]^{+}$ (2)
where $O_{p}=1+m$ and $O_{n}=-m$, with $m$ being a margin hyperparameter. These terms adjust the weight of each sample based on its distance from the ideal similarity value. The margins $\Delta_{p}=1-m$ and $\Delta_{n}=m$ are used to further separate positive and negative pairs in the embedding space.
The logits for positive and negative pairs are then computed as:
$\log_{p}=-\gamma\alpha_{p}(s_{p}-\Delta_{p}),\quad\log_{n}=\gamma\alpha_{n}(s_{n}-\Delta_{n})$ (3)
Finally, the unified Circle Loss function is formulated as follows:
@@
$&\mathcal{L}_{\text{reid}}=\\ &\log\left[1+\sum_{j=1}^{L}\exp\left(\gamma\left(\alpha_{n}^{j}\left(s_{n}^{j}-\Delta_{n}\right)\right)\right)\sum_{i=1}^{K}\exp\left(-\gamma\left(\alpha_{p}^{i}\left(s_{p}^{i}-\Delta_{p}\right)\right)\right)\right]\end{aligned}$
This formulation allows for a dynamic weighting of positive and negative pairs, enhancing the model’s ability to discriminate between similar and dissimilar identities. Circle Loss has proven to be an effective solution for embedding learning in tasks that require fine-grained discrimination, which aligns with the goals of appearance embedding in our tracking framework.
The final total loss is as follows:
$\mathcal{L}_{overall}=\lambda_{reid}*\mathcal{L}_{reid}+\lambda_{det}*\mathcal{L}_{det}$ (5)
where $\lambda_{reid}$ and $\lambda_{det}$ are proportional coefficients used to balance the loss. In this paper, both of them are set to 1. Notably, it should be noted that, when computing the ReID loss, similar to the detection loss, the outputs from each decoder layer are incorporated into the calculation.

### v4

# FastTrackTr: Real-time Multi-Object Tracking with Transformers for Real World

@@ Abstract
Transformer-based multi-object tracking (MOT) methods have attracted significant attention from researchers. However, these Transformer-based models often suffer from suboptimal inference speeds due to their architectural complexities or other inherent issues, rendering them difficult to deploy in practical industrial applications. To address this challenge, we revisited the classic joint detection and tracking (JDT) paradigm and analyzed existing models. Drawing inspiration from DETR’s object queries, which naturally encode object appearance features, we constructed a fast and novel JDT-type MOT framework named FastTrackTr by implementing an efficient inter-frame information transfer mechanism. Benefiting from the superiority of this mechanism, our approach not only reduces the number of queries required for tracking but also avoids introducing excessive network structures, ensuring model simplicity. Experimental results demonstrate that our method not only achieves competitive tracking accuracy across multiple datasets but also enables real-time tracking in practical application environments.
{IEEEkeywords}
Multi-object tracking, Transformers, Real-time
@@ 1 Introduction
[…]
Multi-object tracking is a critical research field in computer vision and industrial intelligence [1, 2, 3, 4, 5], with broad applications in security monitoring, small unmanned ground vehicles, video object analysis, crowd flow detection, and other domains. In recent years, some tracking methodologies based on transformer [6] have garnered significant attention. Meanwhile, these models [7, 8, 9, 10, 11, 12, 13] have also achieved state-of-the-art (SOTA) results on many benchmarks. However, these methods still suffer from a critical issue: their inference speeds remain suboptimal, and the network architectures of some models are highly inconvenient for engineering deployment, making them difficult to apply in practical industrial production. In contrast, some other existing methods [14, 15, 16] exhibit accuracy that may lag behind Transformer-based approaches. Therefore, developing Transformer architectures for real-time tracking remains an urgent research goal.
The limited inference speed of existing methods stems from multiple factors. For query-based tracking models like the MOTR series [9, 10, 11, 12], the primary bottleneck lies not in network architecture design but in hardware acceleration challenges caused by variable query quantities. Notably, computational complexity in many modules scales quadratically with the number of queries. Both our empirical studies and other work [17] confirm that network lightweighting alone yields insufficient speed improvements. The more recent method, MOTIP [13] addresses several MOTR limitations but introduces an auxiliary ID prediction network. Even when substituting MOTIP’s Deformable DETR [18] with RT-DETR [19], inference latency remains unsatisfactory (results shown in Fig. 1). Furthermore, MOTIP’s dynamic parameterization complicates acceleration through TensorRT optimization. Tracking-by-Detection (TBD) paradigms [14, 15], while computationally efficient and deployment-friendly, face their own constraints. When accounting for the combined latency of detection models and re-identification (ReID) networks, only purely kinematic methods [16] achieve competitive speeds—at the cost of inferior performance compared to ReID-enhanced approaches. This raises a fundamental question: Can we develop a framework that simultaneously achieves speed, precision, and practical deployability?
Revisiting historical solutions provides critical insights. The Joint Detection and Tracking (JDT) paradigm was originally designed for rapid tracking but has fallen out of favor in recent years. Nevertheless, its core principles may hold the key to resolving current challenges. Existing transformer-based JDT implementations like TransTrack [7] suffer from limitations similar to MOTIP, while other query-based methods [8, 20, 21] inherit MOTR’s fundamental constraints. CNN-based JDT methods [22, 23], though now less prevalent due to outdated base architectures, demonstrated an effective design: simultaneous detection output and ID embedding generation, followed by association matching. This paradigm inherently preserves inference efficiency while enabling deep coupling of detection features and identity representations through end-to-end joint training. Crucially, JDT aligns naturally with DETR architectures—the object queries produced by DETR inherently encode spatial-semantic information, thereby circumventing the tracking-by-query limitations when repurposed as ID embeddings. Moreover, transformer attention mechanisms surpass CNN counterparts in processing implicit temporal correlations across video frames, potentially enhancing tracking accuracy.
Figure 1: A comparison of the frames per second (FPS) and HOTA of various versions of FastTrackTr and the most advanced transformer-based models on a single NVIDIA RTX 4090. $\text{Dec}^{3}$ refers to a 3-layer decoder, while R18 refers to ResNet18. Additionally, RT-DETR was utilized to construct the MOTIP. The data precision of the scatter plot on the right is PyTorch fp16 and the area of points represents MOTA.
@@
Building upon these insights, Based on the above discussion, we present FastTrackTr, a Fast multi-object Tracking method grounded in the Transformer architecture and aligned with the JDT paradigm. This tracking method achieves impressive tracking progress through a unique decoder that integrates past information. Additionally, due to the ease of generating masks for this method, the batch size can be set relatively large, making our model considerably more training-friendly compared to some of the latest transformer-based MOT methods. Furthermore, the deterministic data architecture inherent in FastTrackTr enables seamless TensorRT acceleration, achieving genuine real-time performance. This computational efficiency remains largely unattainable for contemporary transformer-based MOT methods due to their inherent dynamic computation patterns.
@@
    conducting extensive experiments on multiple benchmark datasets, including DanceTrack [24], SportsMOT [25], MOT17 [26], Visdrone2019 [27], and BDD100k [28], to demonstrate its robust adaptability to various environmental conditions.
  • 4)
    performing physical experiments in industrial settings to confirm its practical feasibility for real manufacturing applications.
@@ 2.1 Transformer-based MOT Methods
The earliest Transformer-based MOT methods were TransTrack [7] and TrackFormer [8]. The former is a JDT model, while the latter introduced the concept of track queries, utilizing these special queries to achieve tracking. Subsequently, MOTR [9] proposed tracklet-aware label assignment (TALA), which optimized label assignment during the training phase to achieve end-to-end tracking. Later improvements to MOTR, such as MOTRv2 [29], MeMOTR [11], MOTRv3 [12], and CO-MOT [10], demonstrated the powerful tracking performance of the MOTR paradigm. Some efforts have been made to address MOTR’s slow inference speed through model lightweighting, but the results have been less than ideal. Recently, alternative Transformer architectures have been proposed. MOTIP’s [13] novel approach of directly predicting tracked object ID numbers is refreshing, while PuTR’s [30] use of Transformers to directly construct a tracking-by-detection paradigm tracker has also been inspirational.
@@ 2.2 Tracking-by-Detection
The TBD paradigm can be considered one of the simplest and most practical tracking methods. Since it essentially associates detected targets, this class of methods often yields satisfactory tracking accuracy as long as the detection results are not poor. The initial SORT algorithm simply utilized the motion state of detected objects for association. Subsequent models like ByteTrack [31] improved tracking accuracy by enhancing detection precision and employing embedding-based Re-identification (Re-id) networks. Some approaches, like OC-SORT [16] and HybridSORT [15], improved upon SORT using purely kinematics-based methods. Later models such as DeepOC-SORT [14] demonstrated that simultaneous improvements in both Re-id network utilization and kinematic models yield the most significant performance enhancements. Currently, there are also algorithms that directly use networks for association.
@@ 2.3 Joint Detection and Tracking
Joint detection and tracking aims to achieve detection and tracking simultaneously in a single stage. This approach was common a few years ago but has been seen less frequently in recent years. Here, we briefly review its development history. JDE [23] and FairMOT [22] learned object detection and appearance embedding tasks from a shared backbone network. CenterTrack [32] localized targets through tracking-conditioned detection and predicted their offsets from the previous frame. MMTrack[5] has constructed a JDT model suitable for multiple scenarios. TransTrack used two decoders to handle tracking and detection separately, then matched and associated the results from both decoders. Our FastTrackTr, by introducing a novel decoder, returns to the essence of the JDT method, simplifies the network structure, and achieves fast tracking.
@@ 3.1 FastTrackTr Architecture
Figure 2: The complete structure of FastTrackTr. For the initial frame, the model functions as a detection system, assigning unique identifiers to each detected object. Subsequently, from the second frame onwards, the historical queries from previous frames and their corresponding masks undergo processing through a single-layer encoder.
The overall structure of FastTrackTr is illustrated in Fig. 2. In frame $T_{1}$, a standard decoder is used to initialize the initial historical information. Subsequently, in the following frames, a Historical Encoder is employed to initialize this historical-representative information. In the next frame, a Historical Decoder, which includes a historical cross-attention layer compared to normal decoder layers, processes this historical information. The tracking principle is very similar to that of JDE and FairMOT, utilizing an additional ID Embedding head to obtain the object’s appearance embedding, which is then used along with detection results for association and matching. This architecture achieves efficient tracking, which can be run in real time on edge devices.
@@ 3.2 Learning Detection and ID Embedding
To enable our FastTrackTr to simultaneously output both the target’s position and id embedding in a single forward pass, ensuring optimal inference speed, we formulate the problem as follows: Given a training dataset $\{(\mathbf{I}_{i},\mathbf{B}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ consisting of image frames $\mathbf{I}_{i}\in\mathbb{R}^{C\times H\times W}$, bounding box annotations $\mathbf{B}_{i}\in\mathbb{R}^{K_{i}\times 4}$, and partial identity labels $\mathbf{y}_{i}\in\mathbb{Z}^{K_{i}}$, we aim to develop a unified model that simultaneously predicts bounding boxes $\tilde{\mathbf{B}}_{i}\in\mathbb{R}^{\tilde{K}_{i}\times 4}$ and discriminative appearance embeddings $\tilde{\mathbf{F}}_{i}\in\mathbb{R}^{\tilde{K}_{i}\times D}$ through a single forward pass.
The model must satisfy two fundamental requirements: (1) accurate bounding box regression through minimization of the localization loss $\mathcal{L}_{\text{det}}(\tilde{\mathbf{B}},\mathbf{B},\tilde{\mathbf{y}},\mathbf{y})$, and (2) temporally consistent feature embeddings where the distance $d_{\phi}(f_{t}^{k},f_{t+\Delta t}^{m})$ between matching identities ($\mathbf{y}_{t}^{k}=\mathbf{y}_{t+\Delta t}^{m}$) is smaller than non-matching pairs ($\mathbf{y}_{t}^{k}\neq\mathbf{y}_{t+\Delta t}^{m^{\prime}}$) by a margin $\alpha>0$. Formally, the joint objective can be expressed as:
$\min_{\theta,\phi}\mathbb{E}\left[\mathcal{L}_{\text{det}}+\lambda\max(0,d_{\phi}(f_{t}^{k},f_{t+\Delta t}^{m})-d_{\phi}(f_{t}^{k},f_{t+\Delta t}^{m^{\prime}}))\right]$ (1)
where $\lambda$ controls the balance between detection accuracy and feature discriminability, with $f_{t}^{k}$ denoting the $k$-th row vector of $\tilde{\mathbf{F}}_{t}$ corresponding to the $k$-th predicted instance.
Our framework builds upon the DETR architecture [18], selected for its inherent compatibility with our technical objectives. The Transformer decoder’s parallel decoding mechanism transforms learnable object queries into interpretable target representations through cross-attention interactions with image features. These joint representations intrinsically encode spatial coordinates and semantic features, which are conventionally mapped to classification probabilities and bounding box coordinates via separate prediction heads.
We extend this paradigm through two strategic modifications: (1) Introducing an auxiliary ID embedding head implemented as a fully-connected layer that projects object queries into a discriminative feature space for target association, and (2) Implementing a joint optimization strategy that preserves detection performance while learning identity-aware representations. The base detection component employs RT-DETR [19] for its real-time efficiency, maintaining its original dual-branch structure and loss formulation:
$\mathcal{L}_{\text{det}}=\underbrace{\mathcal{L}_{\text{box}}(\hat{b},b)}_{\text{Localization}}+\underbrace{\mathcal{L}_{\text{cls}}(\hat{c},c,\text{IoU})}_{\text{Classification}}$ (2)
where $\hat{b},b$ denote predicted/ground-truth boxes and $\hat{c},c$ represent class predictions. The classification loss integrates IoU-aware weighting following RT-DETR, while the box regression combines L1 and generalized IoU losses.
Previous JDT methods [23, 22] suffer from scalability limitations due to their one-hot identity encoding schemes. We instead formulate identity embedding learning as a metric learning problem, decoupling feature dimensions from identity quantities to enhance cross-scene generalizability. This contrasts with recent motion-centric approaches [13, 30] that underutilize spatial features encoded in DETR queries.
@@
Our method employs Circle Loss [33] with adaptive margin control to optimize pairwise similarities in the embedding space. Specifically, by utilizing the dot-product similarity, the id embeddings of the same object across multiple frames are regarded as positive pairs indicated as $s_{p}$, while those between distinct objects are treated as negative samples represented by $s_{n}$. - $\alpha_{p}$ and $\alpha_{n}$ are the adaptive weighting factors for positive and negative samples, determined by the following equations:
$\alpha_{p}$ $=\max(0,O_{p}-s_{p}),\quad\alpha_{n}=\max(0,s_{n}-O_{n})$ (3)
$O_{p}$ $=1+m,\quad O_{n}=-m\quad(m>0)$
where $m$ controls the separation margin. The final loss formulation:
@@
$\mathcal{L}_{\text{reid}}=\log\left[1+\sum_{j}e^{\gamma\alpha_{n}(s_{n}^{j}-\Delta_{n})}\sum_{i}e^{-\gamma\alpha_{p}(s_{p}^{i}-\Delta_{p})}\right]$ (4)
with $\Delta_{p}=1-m$, $\Delta_{n}=m$, and $\gamma$ as the scaling factor. This adaptive weighting mechanism prioritizes hard samples in both positive and negative groups.
The complete objective function combines detection and metric learning losses through layer-wise aggregation across decoder stages:
$\mathcal{L}_{\text{total}}=\lambda_{\text{det}}\mathcal{L}_{\text{det}}+\lambda_{\text{reid}}\mathcal{L}_{\text{reid}}$ (5)
where $\lambda_{\text{det}}=\lambda_{\text{reid}}=1$ ensures balanced optimization. At the same time, in order to better enable the model to implicitly learn temporal information, like other Transformer-based MOT models [9, 13, 8, 7], the loss here is updated after accumulating several consecutive frames.